OpenStudy (anonymous):

Need help .. ? A school system is reducing the amount of dumpster loads of trash removed each week. In week 5, there were 40 dumpster loads of waste removed. In week 10, there were 30 dumpster loads removed. Assume that the reduction in the amount of waste each week is linear. Write an equation in function form to show the amount of trash removed each week. f(x) = - 2x + 40 f(x) = 2x + 40 f(x) = - 2x + 50 f(x) = 2x + 50

OpenStudy (anonymous):

you're given two points : (5, 60) and (10, 40) you need to setup the equation in slope-intercept form : y = mx + b start by finding the slope m first week 5, they dumped 60 week 10, they dumped 40 so it was reduced (5, 60) and (10, 40) x1 y1 x2 y2 slope = y2−y1x2−x1 slope = 40−6010−5 = −205 = −4 substitute m value in above equation y = mx + b y = -4x + b U still need to find B tho use this : the line goes thru (5, 60) so it has to satisfy this point 60=−4(5)+b 60=−20+b 60−(−20)=b 80=b So the answer is c. f(x) = - 4x + 80

OpenStudy (anonymous):

Thank you so much, I understand it a lot better now :)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

you are so welcome welcome